### v1

# Low-Rank Adaptation for Foundation Models: A Comprehensive Review

Abstract
The rapid advancement of foundation models—large-scale neural networks trained on diverse, extensive datasets—has revolutionized artificial intelligence, enabling unprecedented advancements across domains such as natural language processing, computer vision, and scientific discovery. However, the substantial parameter count of these models, often reaching billions or trillions, poses significant challenges in adapting them to specific downstream tasks. Low-Rank Adaptation (LoRA) has emerged as a highly promising approach for mitigating these challenges, offering a parameter-efficient mechanism to fine-tune foundation models with minimal computational overhead. This survey provides the first comprehensive review of LoRA techniques beyond large Language Models to general foundation models, including recent techniques foundations, emerging frontiers and applications of low-rank adaptation across multiple domains. Finally, this survey discusses key challenges and future research directions in theoretical understanding, scalability, and robustness. This survey serves as a valuable resource for researchers and practitioners working with efficient foundation model adaptation.
Index Terms:
Foundation Model, Large Language Models, Low-Rank Adaptation, Parameter-Efficient Fine-Tuning, Multi-Task Learning1 Introduction
Foundation models represent a paradigm shift in artificial intelligence, wherein large-scale neural architectures, pre-trained on extensive and broad datasets, establish generalizable representational frameworks that can be adapted to a wide range of downstream applications [1, 2]. These models span multiple domains, including natural language processing (e.g., GPT-3.5 [3], LLaMA [4]), computer vision (e.g., Swin Transformer [5], MAE [6], SAM [7]), speech processing (e.g., Wav2vec2 [8], Whisper [9]), multi-modal learning (e.g., Stable Diffusion [10], DALL·E 2 [11]), and scientific applications (e.g., AlphaFold [12], ChemBERTa [13], ESM-2 [14]).
Foundation models are characterized by their unprecedented scale, with parameter counts reaching billions or even trillions, and exhibit emergent properties - capabilities that arise spontaneously without explicit training [1]. These architectures have become fundamental building blocks of modern AI systems, enabling breakthrough performance across diverse domains [1, 2] While these models exhibit broad capabilities, task-specific optimization through fine-tuning remains essential for enhancing generalization [15], promoting algorithmic fairness [16], enabling customization [17], and aligning with ethical and societal standards [18, 19]. However, their scale introduces significant computational challenges, particularly in the computational resources required for both training and fine-tuning [20]. Although traditional fine-tuning methods involving full parameters updates have demonstrated effectiveness across various tasks [21, 22], their computational demands often render them impractical for foundation models [23, 24].
Parameter-efficient fine-tuning (PEFT) methodologies have emerged as a solution to these computational challenges [25, 17, 26, 27, 28, 24]. These approaches enable model adaptation by minimizing the number of trainable parameters, substantially reducing computational requirements without compromising task performance. Among these approaches, Low-Rank Adaptation (LoRA) [17] and its variants have gained widespread attention due to their simplicity, empirical effectiveness, and broad applicability across diverse model architectures and domains, as shown in Figure 1.
LoRA is grounded on two key insights that weight updates during fine-tuning often reside in a low-dimensional subspace [29, 30], and task-specific adaptations can be effectively captured using low-rank matrices [17]. By optimizing these low-rank matrices per task while freezing the original model parameters, LoRA achieves efficient adaptation and enables the composition of multiple task-specific adaptations without increasing inference latency [31, 17].
Contributions. This survey provides, to the best of our knowledge, the first comprehensive review of LoRA-based techniques beyond the domain of Large Language Models (LLMs) [32], extending the analysis to the broader landscape of foundation models. Our key contributions are:
(1) Systematic Analysis of Technical Foundations: We provide a structured analysis of the recent technical advances of LoRA, including parameter efficiency strategies, rank adaptation mechanisms, training process improvements, and emerging theoretical perspectives.
(2) Extensive Investigation of Emerging Frontiers: We explore emerging research frontiers, including advanced architectures incorporating multiple LoRA compositions and mixture-of-experts approaches, as well as methodologies for continual learning, unlearning, federated learning, long-sequence modeling, and efficient serving infrastructure.
(3) Comprehensive Review of Applications: We present a comprehensive review of practical applications across diverse domains, including natural language processing, computer vision, speech recognition, scientific discovery, and specialized applications in code engineering, recommender systems, graph learning, and spatial-temporal forecasting.
2 Basics
LoRA [17] constitutes a substantial advancement in parameter-efficient fine-tuning (PEFT). Although originally developed for LLMs, subsequent research has demonstrated its effectiveness across a diverse of foundation models.
The mathematical formulation of LoRA centers on constraining the update matrix to be low-rank during fine-tuning, as shown in Fig. 2, which is implemented through matrix decomposition:
where , , and the rank . By restricting to be low-rank, LoRA minimizes the number of parameters that need to be learned during the fine-tuning process, resulting in significant computational and storage efficiency.
Parameter Initialization Strategies. LoRA employs specific initialization strategies to ensure stable and efficient training. Matrix is typically initialized with values drawn from a random Gaussian distribution, while matrix is initialized with zeros, which ensures that at the start of training, is effectively a zero matrix.
Fine-tuning Process. In LoRA, the fine-tuning process follows these key principles:
The original pretrained weights are kept frozen and do not receive gradient updates during training.
The low-rank matrices and contain the only trainable parameters, capturing task-specific adjustments.
Both and are applied to the input vector separately and their outputs are combined.
The output is scaled by .
The resulting output vectors are summed element-wise:
where is a scaling factor controlling the magnitude of the low-rank update. When optimizing using Adam [33], tuning the scaling factor becomes roughly analogous to adjusting the learning rate [17], provided that the initialization is scaled appropriately. In practice, the value of can be set based on the rank , eliminating the need for extensive hyperparameter tuning.
Advantages of LoRA over full fine-tuning. LoRA offers several key advantages over full fine-tuning when applied to large foundation models:
(1) Parameter Efficiency. LoRA introduces a minimal set of trainable parameters through low-rank decomposition, typically reducing the number of task-specific parameters by several orders of magnitude. This approach is particularly advantageous in resource-constrained environments and multi-task scenarios where multiple adaptations of a base model are required.
(2) Enhanced Training Efficiency. Unlike conventional full fine-tuning, which updates all model parameters, LoRA optimizes only the low-rank adaptation matrices. This approach substantially reduces computational costs and memory requirements, especially for models with billions of parameters. The reduced parameter space typically leads to faster convergence during training.
(3) None-latency Inference. LoRA does not introduce additional inference latency since the update matrix can be explicitly incorporated into the original frozen weights . This integration ensures that the adapted model maintains efficiency during deployment and inference.
(4) Flexible Modular Adaptation. LoRA enables the creation of lightweight, task-specific adapters that can be interchanged without modifying the base model architecture. This modularity facilitates efficient multi-task learning and task switching while minimizing storage requirements compared to maintaining separate model instances for each task.
(5) Robust Knowledge Retention. By preserving the pre-trained weights, LoRA effectively mitigates catastrophic forgetting, a common challenge in conventional fine-tuning. This approach maintains the model’s foundational knowledge while acquiring task-specific capabilities.
(6) Versatile Deployment. The compact nature of LoRA adaptations facilitates efficient deployment and system integration. Multiple adaptations can be readily combined or alternated across different tasks or domains, offering enhanced flexibility compared to traditional fine-tuning approaches.
Through these advantages, LoRA enables efficient adaptation of foundation models while maintaining model performance and significantly reducing computational requirements.
3 Foundations
In this section, we examine the fundamental technical aspects of LoRA across four critical dimensions: parameter efficiency enhancement, rank adaptation strategies, training process refinements, and theoretical foundations. These components constitute the technical foundation of LoRA’s effectiveness.
3.1 Parameter Efficiency Enhancement
Despite the parameter efficiency gains achieved through LoRA with its project-down and project-up matrices, the method still requires a significant number of trainable parameters. For instance, applying LoRA to the LLaMA-2-70B model [4] necessitates updating over 16 million parameters [34], surpassing the total parameter count of some BERT architectures [35]. Current research addresses this challenge through four primary approaches: parameter decomposition, pruning, freezing and sharing, and quantization. Fig. 4 illustrates examples of these techniques.
3.1.1 Parameter Decomposition
Parameter decomposition methods achieve parameter efficiency by decomposing matrices in more compact forms while maintaining task performance. Beyond reducing trainable parameters, these methods also enable more granular control during fine-tuning. Current methodologies can be categorized into two principal approaches: update matrix decomposition [36, 37, 34, 38], and pre-trained weight decomposition [39].
(1) Update Matrix Decomposition. In update matrix decomposition approaches, two primary strategies have emerged: singular value decomposition (SVD) based method and tensor train (TT)-based decomposition.
(i) SVD-based Methods. AdaLoRA [36] parameterizes the updates weights in the form of SVD [40]:
where and represent the left and right singular vectors of , respectively, and the diagonal matrix contains the singular values. AdaLoRA dynamically adjusts the rank of based on importance scoring, enabling adaptive parameter efficiency during fine-tuning. Building upon this, BiLoRA [37] extends this framework with bi-level optimization, separating singular vector and value training across different data subsets to mitigate overfitting.
(ii) TT-based Decomposition. LoRETTA [34] takes a different approach by employing TT decomposition [41], which represents a matrix into a series of low-rank, small, three-dimensional tensors, commonly referred to as cores. Given a matrix , it is first reshaped into a tensor , where . The TT representation of can be formulated as:
where represents a core tensor, and denotes TT rank with . This decomposition reduces the parameter count from to .
LoRETTA introduced two variants: LoRETTA and LoRETTA. LoRETTA employs tensorized adapters, inserting these lightweight modules after each attention and feed-forward sub-layer in the transformer blocks. LoRETTArep, on the other hand, reparameterizes the weight matrix with tensor factors, offering an even more compact PEFT approach. It updates the weights using two unbiased tensorized layers, further reducing the number of trainable parameters while maintaining comparable performance.
TT-LoRA [38] applies this concept directly to the low-rank matrices in the original LoRA formulation. Note that TT-LoRA operates as a parallel adapter, directly modifying the update matrices in the original LoRA formulation. In contrast, LoRETTAadp functions as a series of adapters inserted into the pre-trained model architecture.
(2) Pre-trained Weight Decomposition. DoRA [39] decomposes the pre-trained weight into magnitude and directional components by normalization method:
where is initialized as the magnitude vector , is initialized as and kept frozen, and denotes the vector-wise norm of a matrix across each column. During fine-tuning, the weight is adapted as:
where becomes trainable and represents the LoRA update to the directional component. This decomposition enables independent optimization of magnitude and direction during fine-tuning, leading to learning patterns that more closely resemble full fine-tuning.
Both approaches offer unique advantages in terms of parameter efficiency and fine-tuning flexibility. The update matrix decomposition methods focus on decomposing the incremental updates applied during fine-tuning, while pre-trained weight decomposition directly modifies the structure of the original model weights, where Table I provides a detailed comparison of these methods.
| Method | Strategy | Mechanism | Core Innovation |
| Parameter Pruning | |||
| SparseAdapter [42] | Importance-based | Parameter scoring | Multi-criteria importance evaluation (magnitude, gradient, sensitivity) |
| SoRA [43] | Regularization-based | Gated sparsification | L1-regularized gating for adaptive sparsity |
| LoRA-Drop [44] | Output-based | Layer impact analysis | Dynamic pruning based on layer-wise output contributions |
| Parameter Freezing | |||
| LoRA-FA [45] | Selective freezing | Fixed feature extraction | Random initialization and freezing of matrix |
| Asymmetric LoRA [46] | Theoretical design | Orthogonal projection | Random orthogonal with theoretical guarantees |
| Parameter Sharing | |||
| VeRA [47] | Full sharing | Vector-based adaptation | Shared frozen matrices with trainable scaling vectors |
| NOLA [48] | Basis sharing | Linear combination | Shared basis matrices with trainable coefficients |
| Tied-LoRA [49] | Flexible sharing | Layer-wise tying | Unified framework for cross-layer parameter sharing |
3.1.2 Parameter Pruning
Parameter pruning techniques focus on assessing the importance of different parameters within the LoRA matrices and removing those deemed less important. These methods can be categorized based on their pruning approaches: importance-based pruning, regularization-based pruning, and output-based pruning.
(1) Importance-based Pruning. These methods evaluate parameter importance using multiple metrics. SparseAdapter [42] applies traditional network pruning techniques to LoRA parameters, assessing importance through parameter magnitude, gradient information, and sensitivity analysis. RoseLoRA [50] extends this concept by implementing sensitivity-based scoring for row/column pruning, enabling selective knowledge updates while preserving low-rank adaptation benefits.
(2) Regularization-based Pruning. Regularization-based pruning techniques induce sparsity through optimization constraints. SoRA [43] utilizes a gating mechanism between the down-projection and up-projection matrices of LoRA, utilizing proximal gradient descent with L1 regularization. This approach enables automatic sparsification during training, with zero-valued elements eliminated post-training.
(3) Output-based Pruning. Output-based methods evaluate LoRA parameters based on their layer-wise impact. LoRA-drop [44] evaluates the importance of LoRA modules by analyzing the distribution of across different layers. The method retains individual LoRA modules for the most important layers while sharing a single LoRA across other layers deemed less critical. The importance score computation utilizes , where represents the sampled dataset.
3.1.3 Parameter Freezing and Sharing
Parameter freezing and sharing techniques reduce trainable parameters through matrix-wise Freezing and cross-layer parameter sharing.
(1) Matrix-wise Freezing. Research has revealed asymmetric roles of matrices and in adaptation. LoRA-FA [45] demonstrates that freezing a randomly initialized matrix while only updating can maintain model performance. Asymmetric LoRA[46] provides theoretical foundations for this approach, showing that primarily acts as a feature extractor while serves as a task-specific projector. This leads to an enhanced design using a frozen random orthogonal matrix for , further reducing parameters while preserving performance.
(2) Cross-layer Parameter Sharing. Several methods explore parameter sharing across network layers. VeRA [47] shares frozen matrices and across layers, training only scaling vectors for adaptation. NOLA [48] extends this concept by representing and as trainable linear combinations of shared frozen basis matrices. Tied-LoRA [49] implements layer-wise parameter tying while keeping the shared matrices trainable, offering a flexible framework that includes VeRA as a special case when the shared matrices are frozen.
3.1.4 Parameter Quantization
Quantization [51, 52, 53] optimizes neural network complexity through lower-precision numerical representations, substantially reducing storage and computational requirements. For a comprehensive quantization background, readers may refer to [54]. In LoRA contexts, quantization approaches are characterized by two primary dimensions: quantization timing and quantization techniques.
(1) Quantization Timing. Quantization timing refers to when quantization occurs before, during, or after fine-tuning.
Pre-finetuning quantization. Pre-finetuning quantization is that the pretrained weights are quantized prior to any LoRA-based adaptation. For example, QLoRA [55] employs a 4-bit NormalFloat (NF4) quantization method. Similarly, LoftQ [56] improves upon this by addressing discrepancies introduced by quantizing high-precision weights. LoftQ jointly quantizes the pretrained model while optimizing low-rank initializations using an iterative algorithm.
During-finetuning quantization. During-finetuning quantization applies quantization both before and throughout the fine-tuning process. Methods like QA-LoRA [57] leverage group-wise quantization to adjust the precision dynamically during training, ensuring a more balanced interaction between low-rank updates and quantized weights.
Post-finetuning quantization. Post-finetuning quantization, such as in LQER [58], occurs after the fine-tuning is completed, focusing on quantization primarily for inference. LQER utilizes a low-rank SVD-based decomposition to minimize quantization errors, ensuring that the quantized weights closely match the original high-precision weights.
| Method | Timing | Target | Precision | Technique | Low-Rank | Optimization | Memory Focus | Dequant |
| QLoRA [55] | Pre-FT | Pretrained | 4 bit | NormalFloat | Standard LoRA | Separate | FT&Inference | Partial |
| QA-LoRA [57] | Pre & During FT | Pretrained | 2, 3, 4 bit | Group-wise | Q-aware LoRA | Joint | FT&Inference | None |
| LoftQ [56] | Pre-FT | Pretrained | mixed | Uniform & NormalFloat | Q-aware LoRA | Joint | FT | Partial |
| LQER [58] | Post-FT | Pretrained | mixed | Group-wise & adaptive | SVD-based LR | Q-error min | Inference | None |
| QDyLoRA | During FT | Pretrained | mixed | Rank sampling | Dynamic LoRA | Rank selection | FT | None |
| LQ-LoRA [59] | Pre-FT | Pretrained | mixed | ILP & data-aware | Q-aware LoRA | Joint | FT&Inference | Partial |
FT = Fine-tuning, Q = Quantization, LR = Low-Rank, ILP = Integer Linear Programming
(2) Quantization Technique. Different approaches include uniform quantization, non-uniform quantization, and mixed-precision quantization have been proposed for LoRA.
Uniform Quantization. Uniform quantization assigns the same bit-width across all weights, regardless of their distribution. QA-LoRA [57] applies uniform quantization with group-wise refinement, which balances the precision trade-offs to optimize memory efficiency and adaptation. However, uniform quantization may struggle with non-uniformly distributed weights, which is where non-uniform quantization becomes more effective.
Non-uniform Quantization. QLoRA [55] employs non-uniform quantization, designed specifically for weights with Gaussian-like distributions, which helps allocate more precision where it is needed most, near zero. This approach allows for a better representation of smaller weights, which dominate in pretrained models.
Mixed-precision Quantization. Mixed-precision quantization offers even more flexibility by dynamically adjusting the bit-width based on the weight matrix or layer. Methods such as LoftQ [56] and LQ-LoRA [59] leverage mixed precision to optimize quantization across different components of the model. LoftQ, for example, alternates between quantizing the residuals of the weight matrix and refining the low-rank components using SVD. By iteratively optimizing the low-rank parameters and adjusting the quantization levels, LoftQ is able to minimize quantization errors. LQ-LoRA expands upon this approach by employing integer linear programming to dynamically configure the bit-width for each weight matrix, and also introduces a data-aware mechanism that uses an approximation of the Fisher information matrix to inform the quantization process. This allows LQ-LoRA to achieve a more accurate decomposition of the weight matrices with minimal quantization-induced loss.
To summarize, pre-finetuning quantization approaches, such as QLoRA and LoftQ, generally provide greater memory savings by freezing the pretrained weights, while post-finetuning methods like LQER focus more on refining precision for inference. In terms of quantization techniques, non-uniform and mixed precision methods, as seen in QLoRA, LoftQ, and LQ-LoRA, demonstrate superior performance in low-bit scenarios by offering more flexible precision allocation based on weight distribution. The timing of quantization and the specific quantization technique both play critical roles in determining the balance between memory efficiency and model performance. Table III provides a comprehensive summary of the discussed quantization methods.
Overall, parameter efficiency enhancement techniques in LoRA has evolved through four complementary approaches: decomposition, pruning, freezing and sharing, and quantization, each offering unique trade-offs between model performance and resource utilization. Building upon these efficiency gains, we next explore how rank adaptation strategies further advance LoRA’s capabilities.
3.2 Ranking Adaptation
Rank is a crucial parameter in LoRA, directly impacting the model adaptability and number of trainable parameters. The original LoRA method employs a fixed low rank across all layers, which may not be optimal for different downstream tasks and model architectures. To address limitations, recent works have proposed various approaches to optimize rank allocation in LoRA, which can be broadly categorized into two main aspects: rank refinement and rank augmentation. Figure 5 presents an illustration of these two methods.
3.2.1 Rank Refinement
Rank refinement methods aim to adaptively select the rank of LoRA modules during the fine-tuning. The key insight is that different layers may require varying degrees of adaptation and thus benefit from different ranks. Rank refinement approaches can be grouped into three main types: adaptive allocation, heuristic strategies, and multi-rank training.
(1) Adaptive Allocation. Adaptive allocation methods dynamically adjust the ranks of LoRA modules during training based on importance metrics derived from the data or model parameters. AdaLoRA [36] introduces an adaptive mechanism for rank allocation by parameterizing LoRA updates using SVD. It dynamically prunes singular values based on their magnitudes, allowing each layer to have a tailored rank while maintaining a global parameter budget. Similarly, SoRA [43] employs a learnable gating mechanism to control the effective rank of each LoRA module. The gates are optimized using proximal gradient descent with regularization to promote sparsity. This approach enables the automatic discovery of suitable ranks for different layers, improving parameter efficiency without manual tuning.
(2) Heuristic Strategies. Heuristic strategies allocate ranks based on predefined rules, which can be from prior knowledge or empirical observations. PRILoRA [60] proposes a deterministic strategy where the rank of LoRA modules increases linearly from lower to higher layers. Motivated by the observation that higher layers often require more adaptation in transfer learning, this heuristic allocates higher ranks to upper layers.
(3) Multi-Rank Training. Multi-rank training methods enable the model to perform well across a range of ranks, offering flexibility during inference. DyLoRA [61] trains LoRA modules across a spectrum of ranks simultaneously. In each training iteration, it samples ranks from a predefined distribution, allowing the model to learn to perform effectively across multiple ranks. This strategy enables adaptability during inference without requiring additional training, which is beneficial in deployment scenarios with varying computational constraints.
3.2.2 Rank Augmentation
Rank augmentation methods aim to achieve high-rank model updates through sequences of low-rank modifications, bridging the performance gap between LoRA and full-parameter fine-tuning. These methods can be categorized into two types: matrix merging-based methods and matrix resampling-based methods.
(1) Matrix merging-based methods. Matrix merging-based methods increase the rank by merging low-rank update matrices. The key idea is that the sum of multiple low-rank matrices can approximate a higher-rank matrix, thereby enhancing the ability to capture complex patterns without incurring substantial computational overhead.
ReLoRA [62] introduces an iterative training framework where low-rank LoRA modules are trained and periodically merged into the pre-trained model weights. By resetting the optimizer and initializing new LoRA modules after each merge, ReLoRA effectively increases the overall rank while maintaining memory efficiency.
COLA [63] proposes a similar iterative optimization strategy inspired by the Frank-Wolfe algorithm [64]. It iteratively trains LoRA modules and merges them into the model, incrementally building a higher-rank adaptation. Each new LoRA module minimizes the residual error from previous adaptations, enabling COLA to achieve high-rank expressiveness without increasing the computational cost per iteration.
MELoRA [65] introduces a parallelization approach to rank augmentation. The core idea is to train multiple small LoRA modules concurrently and concatenate their outputs to form a higher-rank adaptation. By assembling a mini-ensemble of low-rank adapters, MELoRA constructs an equivalent block-diagonal matrix that collectively has a higher rank.
XGBLoRA [66] introduces the Gradient Boosting (GB) framework for LoRA. It assembles the final model by combining a sequence of Rank-1 boosters (LoRA adaptations), progressively refining the model’s predictions. With the GB principle of the weak learner (i.e., strong ensemble model from a set of weak predictors), XGBLoRA overcomes the dilemma between extreme low-rank adaption and effectiveness.
(2) Resampling-based method. Matrix resampling-based methods achieve high-rank adaptations by dynamically resampling the projection matrices during training. The fundamental idea is to utilize time to accumulate the effect of high-rank updates while operating with low-rank matrices at each training step.
FLoRA [67] reinterprets LoRA as a gradient compression and decompression mechanism. The motivate that FLoRA periodically resamples the projection matrices used in LoRA modules during training. By changing these matrices at scheduled intervals, the method ensures that different subspaces are explored over time, effectively accumulating a higher-rank adaptation.
| Adapter | Adjustment | Method |
| AdaLoRA[36] | Refinement | Adaptive Allocation |
| SoRA[43] | Refinement | Adaptive Allocation |
| PRILoRA[60] | Refinement | Heuristic Strategies |
| DyLoRA[61] | Refinement | Multi-Rank Training |
| ReLoRA[62] | Augmentation | Matrix Merging-Based |
| COLA[63] | Augmentation | Matrix Merging-Based |
| MELoRA[65] | Augmentation | Matrix Merging-Based |
| XGBLoRA[66] | Augmentation | Matrix Merging-Based |
| FLoRA[67] | Augmentation | Resampling-Based |
In summary, rank adaptation strategies enhance LoRA adaptability by tailoring the rank of adaptation matrices to better suit the requirements of different layers and tasks. Table IV presents a detailed summary about rank refinement and augmentation.
3.3 Training Process Improvements
While LoRA has demonstrated remarkable success in parameter-efficient fine-tuning, optimizing its training dynamics remains crucial for maximizing adaptation performance. In this section, we discuss recent advancements aimed at improving the training process, especially learning rates, dropout strategies, and scaling factors.
Learning Rate. In standard LoRA fine-tuning, a uniform learning rate is typically applied to both low-rank matrices and . However, Hayou et al. [68] observe that this practice leads to suboptimal performance, especially as the model width increases. The issue lies in the updates to and contribute differently to the learning dynamics. To address this limitation, Hayou et al. [68] propose LoRA+, a method that assigns different learning rates to matrices and . Their theoretical analysis in the infinite-width limit reveals that for efficient learning, the magnitudes of feature updates from both and should be . This necessitates scaling the learning rates such that and , where denotes the model width. In practice, LoRA+ introduces a fixed ratio , allowing practitioners to tune a single learning rate while automatically adjusting the other.
Dropout Strategies. Despite the reduced number of trainable parameters in LoRA-based models, overfitting remains a concern, particularly when fine-tuning small or specialized datasets. Traditional dropout techniques may not suffice to mitigate overfitting in this context. Wang et al. [69] highlight this vulnerability and propose a comprehensive framework to address it through dropout along three dimensions: dropping position, structural pattern, and compensation measure. The dropping position specifies where the noise is introduced, such as in the attention logits, weights, or hidden representations. The structural pattern defines the granularity of unit deactivation, encompassing element-wise, column-wise, or span-wise patterns. The compensation measure aims to minimize the discrepancy between training and inference phases by incorporating techniques like normalized rescaling or Kullback-Leibler divergence loss. Building on this framework, the authors present HiddenKey [69], a dropout method that combines column-wise dropout of attention logits with element-wise dropout of hidden representations, supplemented by a KL divergence loss.
Scaling Factor. In LoRA, a scaling factor is applied. However, as Kalajdzievski [70] points out, this scaling factor can cause gradient collapse when increasing the adapter rank, resulting in slowed learning and diminished performance for higher-rank adapters. To overcome this limitation, Kalajdzievski [70] proposes rsLoRA, which redefines the scaling factor to be . This adjustment ensures that the adapters are rank-stabilized, meaning that both the forward and backward pass maintain stable magnitudes relative to the rank, even as it becomes large. Theoretically derived in the infinite-width limit, this scaling factor prevents gradient collapse, enabling stable learning across different adapter ranks.
By adaptive learning rates to the distinct roles of LoRA matrices, mitigating overfitting through structured dropout, and preventing gradient collapse with rank-stabilized scaling, these methods enhance both the efficiency and effectiveness of LoRA fine-tuning. We next examine the theoretical foundations underlying LoRA’s performance.
3.4 Theoretical Foundations
While the practical advantages of LoRA are evident, understanding its underlying principles from a theoretical perspective is crucial. This section addresses key questions regarding its effectiveness, optimal rank selection, roles of update matrices, and induced behavioral changes in theoretical aspects.
Q1: Why does LoRA work effectively? LoRA achieves competitive performance with full fine-tuning while updating only a small subset of parameters. This phenomenon can be understood through the Neural Tangent Kernel (NTK) theory111NTK describes the evolution of neural network gradients during training. Malladi et al. [71] show that LoRA approximately preserves the kernel of the original model during fine-tuning. Specifically, with high probability [71],
where and are the kernels induced by LoRA and full fine-tuning respectively, is the number of examples in the dataset, is an upper bound on the L2 norms of gradients and inputs, is the approximation error, and is the probability bound given by , where is the rank used in LoRA.
Although LoRA restricts updates to a low-rank subspace, it effectively targets the gradients most responsible for significant transformations in network behavior based on the Equation (7). By focusing on these critical gradients, LoRA preserves the model’s ability to generalize, ensuring that the network remains sensitive to essential input variations while being highly parameter-efficient.
Q2: How many ranks are required for optimal LoRA performance? The rank in LoRA fine-tuning is crucial for understanding the expressivity of adaptation and maintaining computational efficiency.
Zeng and Lee [72] conducted a comprehensive study on the expressive power of LoRA across different architectures. (i) For fully connected neural networks, that LoRA can adapt any model to accurately represent a smaller target model if the LoRA-rank satisfies:
(ii) For Transformer networks, they demonstrate that any model can be adapted to a target model of the same size with rank- LoRA adapters. These findings provide a theoretical foundation for determining the minimum rank necessary for effective adaptation across different architectures.
Complementing this work, Jang et al. [73] analyzed LoRA training in the NTK regime, yielding several key insights: (i) They proved that full fine-tuning (without LoRA) admits a low-rank solution of rank , where is the number of training data points. (ii) Using LoRA with rank eliminates spurious local minima, facilitating efficient global minima discovery. This result suggests a lower bound for the LoRA rank to ensure optimization stability. (iii) They provided generalization guarantees for LoRA-adapted models, demonstrating that the generalization error is bounded by . This bound offers reassurance about the performance of LoRA-adapted models on unseen data.
These theoretical analyses offer valuable guidance for hyperparameter tuning in LoRA applications.
Q3: What are the roles of update matrices and B? Zhu et al. [46] provide a comprehensive analysis of the distinct roles played by matrices and in LoRA. Their work reveals an inherent asymmetry in these matrices, which has important implications for fine-tuning efficiency and model generalization.
The authors [46] demonstrate that primarily functions as a feature extractor from the input, while projects these features towards the desired output. This asymmetry suggests that fine-tuning alone can be more effective than fine-tuning . Notably, their analysis shows that a randomly initialized can perform nearly as well as a fine-tuned one, challenging the conventional practice of updating both matrices. Building on this insight, Zhu et al. derive generalization bounds for different LoRA variants using an information-theoretic framework. When fine-tuning only B, the generalization bound takes the form:
where is the rank, is the quantization bits, relates to the sub-Gaussianity of the loss, is the sample size, and is the output dimension of the -th layer. This bound is tighter compared to updating both and , suggesting that freezing as a random orthogonal matrix and only updating could potentially enhance the generalization to unseen data.
These findings align with and extend the insights from previous questions, particularly the discussion on optimal rank selection. By focusing on updating alone, researchers can potentially achieve better generalization while further reducing the number of trainable parameters, thus enhancing both the efficiency and effectiveness of LoRA fine-tuning.
Q4: What behavioral changes does LoRA induce in the model? Koubbi et al. [74] analyzed the dynamics of attention matrices, demonstrating that LoRA-induced low-rank modifications maintain short-term stability in token clustering while facilitating significant long-term divergence in learned representations.
LoRA updates attention matrices with low-rank matrices as , introducing controlled perturbations:
Token dynamics under LoRA are described by:
where attention weights are based on the softmax of the Query and Key matrices.
LoRA maintains short-term stability of token clustering, with the Wasserstein distance between perturbed and unperturbed token distributions remaining bounded:
A key result is the identification of a phase transition, where tokens bifurcate into new clusters after a critical time , governed by the eigenvalue gap of the Value matrix. This shows how LoRA fine-tunes models without catastrophic forgetting, preserving token structure early in training while allowing controlled divergence.
These theoretical foundations of LoRA show its effectiveness, from its competitive performance explained by the NTK theory to its ability to prevent catastrophic forgetting through controlled token dynamics. The insights into optimal rank selection and the asymmetry of update matrices offer practical guidelines for improvements.
4 Frontiers
Building upon the technical foundations discussed above, which establish the core components and mechanisms of LoRA, this section explores frontier developments that extend the capabilities of LoRA in novel directions. These frontier developments leverage and combine its fundamental principles to enable new functionalities, tackle more complex tasks, and address challenges in model adaptation.
4.1 Advanced Architecture
While the original LoRA method significantly enhanced the efficiency of fine-tuning and demonstrated performance comparable to full fine-tuning, it had limitations in flexibility, generalization, and handling multiple diverse tasks simultaneously. To address these limitations, researchers have developed advanced LoRA architectures to further improve performance, parameter efficiency, and generalization ability.
4.1.1 LoRA Composition
One major innovation in advanced LoRA architectures is the dynamic composition of multiple LoRA modules to enhance adaptability and generalization across diverse tasks.
Optimization-based Composition. LoRAHub [75] leverages CMA-ES [76] gradient-free optimization to determine optimal coefficients for combining LoRA modules. Through few-shot learning, it autonomously selects and integrates modules for new tasks without requiring manual expertise or gradient computation. Similarly, LoRA-Flow[77] introduces dynamic fusion weights to adjust the impact of different LoRAs at each generation step, determined by a fusion gate with minimal parameters. This approach outperforms baselines with static task-level fusion weights across various
Retrieval-based Composition. LoraRetriever [78] implements dynamic retrieval and composition of LoRA modules based on input prompts. It first embeds task-specific LoRAs into a shared space using instruction fine-tuning on a subset of tasks, then retrieves relevant modules using cosine similarity. The framework supports both module fusion and mixture strategies while maintaining efficient batch processing.
Batch-oriented Composition. FLORA [79] enables each example in a minibatch to utilize unique low-rank adaptation weights through efficient matrix operations. This design significantly improves throughput and reduces latency compared to traditional batched approaches, particularly beneficial when serving diverse user requests in production environments.
By enabling models to select and combine multiple LoRA modules based on the task or input, these methods overcome the limitations of standard LoRA in handling diverse tasks and improve overall performance.
4.1.2 Generalized Framework
Another advancement involves extending the LoRA architecture itself to capture both task-specific and general features more effectively.
Dual-branch Framework. Hydra [80] presents a more generalized formulation by integrating both parallel and sequential LoRA branches within the model. The parallel branch learns task-specific features, similar to standard LoRA, while the sequential branch linearly combines pre-trained features. This dual branch enables Hydra to capture both task-specific adaptations and leverage general pre-trained knowledge, offering a comprehensive adaptation mechanism that improves performance across tasks.
Multi-PEFT United Framework. GLoRA [81] further make a generalization by unifying various parameter-efficient fine-tuning methods beyond LoRA. It introduces trainable support tensors to scale and shift weights, features, and biases, effectively subsuming methods like LoRA, adapter tuning, and prompt tuning within a single framework. GLoRA employs evolutionary search to determine optimal layer-wise configurations of these tensors, which can take scalar, vector, or low-rank matrix forms. Through structural re-parameterization, GLoRA incurs no additional inference cost while providing greater flexibility than previous PEFT methods.
These generalized architectures enhance the expressive power of LoRA by incorporating additional mechanisms for capturing diverse features and facilitating more effective fine-tuning across tasks.
| Method | Key Features and Mechanism |
| LoRAHub [75] | Optimization-based composition using CMA-ES for gradient-free weight optimization; enables few-shot adaptation without manual expertise. |
| LoraRetriever [78] | Retrieval-based composition; supports both fusion and mixture strategies with efficient batch processing. |
| LoRA-Flow [77] | Dynamic fusion with fusion gates to adjust LoRA impact in generative tasks. |
| FLORA [79] | Batch-oriented composition; achieves 3x throughput improvement at low ranks. |
| Hydra [80] | Dual Branch; Combines parallel branch (new features) and sequential branch (pre-trained features). |
| GLoRA [81] | Multi-PEFT unified framework with trainable tensors for weights/features/bias. |
| XGBLoRA [66] | Gradient boosting framework with rank-1 LoRA adapters; combinations of multi-weak learners. |
4.1.3 Gradient Boosting with LoRA
Gradient Boosting with LoRA (GBLoRA) combines weak learners through iterative LoRA module training to minimize residual errors. After boosting iterations, the finetuned model is expressed as:
with cumulative model updates:
where controls the contribution of each LoRA booster . The weak learner principle enables GBLoRA to achieve strong performance with low-rank updates. XGBLoRA [66] established convergence guarantees and expressiveness bounds, demonstrating how increased boosting iterations can compensate for lower ranks. This framework unifies various matrix merging methods like ReLoRA [62], COLA [63], and MeLoRA [65] within the GB paradigm.
4.1.4 Mixture of Experts with LoRA
Another important branch in the development of advanced LoRA architectures is the combination of LoRA with Mixture of Experts (MoE). MoE is a neural network architecture where multiple “expert” sub-networks specialize in different input patterns [82]. A gating mechanism routes inputs to the most appropriate experts, allowing the model to handle a wide range of tasks efficiently [83]. Given the input , the MoE model computes
where is the output, is a gating function, is an expert, and is the number of experts.
By integrating LoRA with MoE, models learn multiple pairs of low-rank matrices (LoRA experts) instead of a single pair, with a router determining the weights or selection of experts based on the inputs. During fine-tuning, the pre-trained LLM weights remain fixed, while the LoRA experts and the router are trained, leveraging the parameter efficiency of LoRA and the specialization capabilities of MoE. The typical framework is illustrated in Figure 6 (B).
Research on LoRA-MoE methods can be broadly categorized into three groups based on their primary objectives: (1) enhancing performance and parameter efficiency, (2) preserving knowledge during fine-tuning, and (3) adapting to multi-task learning. While these categories highlight different focuses, many approaches address multiple objectives simultaneously.
(1) Efficiency-oriented Design. Methods in this category aim to match full fine-tuning performance with minimal parameter overhead.
Zadouri et al. [84] introduced MoV and MoLoRA, aiming to achieve full fine-tuning performance while updating less than 1% of parameters and improving generalization to unseen tasks. MoV and MoLoRA utilize (IA)3 vectors and LoRA adapters as experts, respectively, employing a soft merging strategy where all experts contribute to the output weighted by router probabilities.
Building on these concepts, Luo et al.[85] proposed MoELoRA, treating LoRA modules as experts within an MoE framework. MoELoRA incorporates multiple LoRA experts and a gating network, employing top- routing and a load-balancing loss[86] to prevent convergence towards a limited set of experts. The application of contrastive learning among experts mitigates the random routing problem often observed in MoE models [87].
However, the fixed number of LoRA experts, e.g., MoELoRA [85], lacks flexibility and can be redundant due to representation collapse or learned routing policy overfitting [88]. To address this, Gao et al. [89] introduced MoLA, a layer-wise expert distribution approach that enables flexible assignment of LoRA experts across different Transformer layers. MoLA employs a top- routing mechanism to select relevant experts for each input. Beyond improving performance and parameter efficiency, MoLA exhibited promising continual learning capabilities due to its sparse expert activation, allowing the model to retain knowledge from previous domains while adapting to new ones.
(2) Memory-based Adaptation. These approaches focus on preventing catastrophic forgetting during adaptation. Two notable approaches, LoRAMoE [90] and MoRAL [91], address the challenge of knowledge retention while adapting LLMs to new tasks or domains.
LoRAMoE [90] introduces multiple LoRA experts integrated via a router network, using a localized balancing constraint to encourage some experts to focus on leveraging world knowledge for downstream tasks. It employs a top- routing strategy, enabling the model to maintain world knowledge while improving performance on multiple tasks. MoRAL [91] uses question-answer pairs from unstructured text and combines the multi-tasking abilities of MoE with LoRA’s parameter efficiency. It employs a soft routing mechanism where all experts contribute to the output weighted by router probabilities. MoRAL maintains performance on previously seen tasks while adapting to new domains, addressing catastrophic forgetting.
| Method | Routing | Key Feature |
| MoV/MoLoRA [84] | Soft routing | All experts contribute with weights |
| MoELoRA [85] | Top-k | Contrastive between experts |
| MoLA [89] | Top-k | Layer-wise expert distribution |
| LoRAMoE [90] | Top-k | Localized balancing for knowledge |
| MoRAL [91] | Soft routing | Lifelong learning framework |
| MOELoRA [92] | Task-based | Task identifier conditioning |
| MoCLE [93] | Cluster-based | Instruction cluster routing |
| LLaVA-MoLE [94] | Token-level | Top-1 sparse expert selection |
(3) Task-based Integration These methods tackle domain specificity and task interference challenges. Domain specificity arises when models trained on general-purpose data lack the specialized knowledge required for specific domains like medicine [92] or finance [95]. Task interference occurs when multiple tasks and their datasets compete during training, leading to degraded performance across tasks [94, 93, 95].
To address domain specificity, Liu et al. [92] proposed MOELoRA for multi-task medical applications. MOELoRA introduces multiple LoRA experts, each consisting of low-rank matrices, with a task-motivated gating function controlling each expert’s contribution based on task identities. This approach allows for task-specific learning while maintaining a shared knowledge base across tasks. Feng et al. [95] introduced MOA, an end-to-end parameter-efficient tuning method for multi-task learning. MOA first trains individual LoRA modules for different tasks, and then combines them using a sequence-level routing mechanism based on domain metadata, allowing the flexible combination of domain-specific LoRAs. Buehler et al. [96] presented X-LoRA, employing a dynamic MoE strategy with a deep layer-wise token-level approach. Starting with pre-trained LoRA adapters, X-LoRA uses a gating strategy leveraging hidden states to dynamically mix adapted layers. This allows the model to create novel combinations to solve tasks, demonstrating strong performance in scientific applications.
To solve task interference, Gou et al. [93] addressed task conflicts in vision-language instruction tuning with MoCLE. This method introduces an MoE architecture that activates task-customized parameters based on instruction clusters, employing a cluster-conditional routing strategy and incorporating a universal expert to improve generalization to novel instructions. Chen et al. [94] proposed LLaVA-MoLE to mitigate data conflicts in instruction fine-tuning of Multimodal LLMs. It introduces a sparse MoE design with multiple LoRA experts and employs a token-level routing strategy where each token is routed to the top-1 expert. This allows adaptive choices for tokens from different domains, effectively addressing data conflicts.
In addition, Tian et al. [97] proposed HydraLoRA, an asymmetric LoRA architecture that challenges the conventional symmetric expert structure in MoE-based approaches. Through empirical analysis, they discovered that in multi-task settings, matrix parameters from different LoRA heads tend to converge while matrix parameters remain distinct. Building on this observation, HydraLoRA introduces an architecture with a shared matrix across all tasks and multiple task-specific matrices, employing a trainable MoE router to automatically identify intrinsic components within the training data.
By employing various routing strategies and expert designs, these methods enable efficient adaptation to multiple tasks or domains while mitigating interference and maintaining task-specific performance. The integration of MoE with LoRA has demonstrated promising results in enhancing performance, preserving knowledge, and facilitating multi-task adaptation across various domains.
4.2 LoRA for Continual Learning
The parameter-efficient nature of LoRA allows for incrementally updating models on new tasks while mitigating catastrophic forgetting [98, 99]. Several key advantages motivate the use of LoRA for Continual Learning (CL): (1) reduced computational costs compared to full fine-tuning, (2) natural isolation of task-specific knowledge, and (3) flexible combination of task-specific adaptations. Existing LoRA-based continual learning methods can be broadly categorized into three approaches: regularization-based method, task arithmetic-based method, and ensemble-based techniques.
Regularization-based approaches that leverage parameter constraints on LoRA updates as the primary mechanism to prevent catastrophic forgetting, focusing on preserving critical model parameters. O-LoRA [98] addresses catastrophic forgetting by constraining new task updates to be orthogonal to the subspace of previous tasks. It leverages the insight that LoRA parameters effectively capture the gradient subspace of a task. O-LoRA incrementally learns new tasks in orthogonal subspaces while keeping previous LoRA parameters fixed. This approach allows for effective knowledge accumulation without interference. Online-LoRA [100] is a task-free online continual learning framework for Vision Transformers that addresses catastrophic forgetting without relying on rehearsal buffers. It combines weight regularization to protect important parameters and loss dynamics monitoring to detect distribution shifts, enabling real-time model adaptation while maintaining performance across changing data streams. It does not assume any task boundaries.
Task arithmetic-based approaches that leverage task vector arithmetic on LoRA parameters [101]. Chitale et.al [101] apply arithmetic operations on LoRA parameters to combine knowledge from multiple tasks. This approach trains separate LoRA modules for each task and then uses task vector addition to create a task-agnostic model. A key insight is that LoRA parameters create semantic ”task vectors” in weight space that can be manipulated algebraically.
Ensemble-based works maintain and combine multiple task-specific LoRA modules. CoLoR [99] maintain separate LoRA modules for each task and use an unsupervised approach to select the appropriate module at inference time. CoLoR trains task-specific LoRA modules sequentially and combines them using prototype-based task identification. This allows for the isolation of task knowledge while enabling flexible combinations. AM-LoRA [102] uses multiple task-specific LoRA modules combined with an attention mechanism to integrate knowledge from different tasks. The attention-based mixing strategy enables adaptive knowledge integration while preventing catastrophic forgetting between tasks.
While these approaches demonstrate the potential of LoRA for continual learning, several challenges remain. The orthogonality constraint in O-LoRA may be overly restrictive for tasks with overlapping knowledge. Task arithmetic assumes tasks can be linearly combined, which may not hold for all scenarios. Ensemble methods face challenges in task identification and scaling to many tasks.
4.3 LoRA for Unlearning
LoRA facilitates the targeted removal of specific knowledge from foundation models without necessitating extensive retraining. This section categorizes and examines methodologies employing LoRA for unlearning, focusing on three primary categories: modular decomposition methods, optimization-based approaches, and progressive unlearning pipeline.
Modular Decomposition Methods. These methods focus on decomposing and modularizing model components to support unlearning. Gao et al. [103] introduce an orthogonal LoRA mechanism that ensures parameter disentanglement during successive unlearning processes. This design ensures that unlearning requests can be processed consecutively without causing interference with retained knowledge. Chen and Yang [104] propose adding efficient unlearning LoRA layers, employing a selective teacher-student objective to guide the model in ”forgetting” specific data. Additionally, Lizzo and Heck [105] introduce UNLEARN, where LoRA layers are adapted to identify and isolate targeted knowledge in low-dimensional subspaces.
Optimization-Based Approaches. These approaches that primarily rely on optimization methods to target and remove specific knowledge. Cha et al. [106] propose a LoRA initialization technique that adapts the low-rank matrices based on the data, weighted by Fisher information to prioritize parameter adjustments crucial to removing targeted knowledge. In a different approach, Gundavarapu et al. [107] explore a gradient-based fine-tuning approach with LoRA to selectively unlearn harmful or unwanted information. Utilizing gradient ascent combined with low-rank LoRA updates, this method refines specific parts of the model’s knowledge.
Sequential Pipeline Strategies. These approaches implement structured, multi-step procedures for systematic unlearning. Liu et al. [108] leverage LoRA to negate specific harmful knowledge in a structured two-stage process. The first stage focuses on identifying harmful content, while the second stage applies LoRA to suppress and neutralize such knowledge without affecting other learned information. This methodical approach ensures systematic removal of unwanted information while preserving the model’s general capabilities.
4.4 LoRA for Federated Learning
In an era of heightened data privacy concerns, Federated Learning (FL) offers a promising approach to leverage collective knowledge while maintaining robust protection of individual data. The integration of LoRA into Federated Foundation Models (FFM) has made foundation models more accessible to resource-constrained devices, particularly in edge computing scenarios, potentially revolutionizing IoT and mobile applications. The combination of federated instruction tuning and value alignment with LoRA creates a powerful synergy that addresses several critical challenges in distributed machine learning.
Privacy and Security. Privacy and security and protection is paramount in federated learning. FedIT [109] and FFA-LoRA [110] establish fundamental frameworks that combine federated learning with instruction tuning for LLMs. These frameworks implement FedAvg privacy-preserving mechanisms, ensuring instruction data remains on local devices while only encrypted LoRA parameters are transmitted and aggregated at the central server. Taking a different approach, PrivateLoRA [111] transmits only activations between the central cloud and edge devices to ensure data locality. Advancing security further, Huang et al. [112] integrate model slicing with Trusted Execution Environment (TEE) to guard against malicious attacks, employing server-side TEE for latter model layers and Sparsification Parameter Fine-tuning (SPF) with LoRA, achieving both security and performance without requiring client-side TEE.
Computation Efficiency. Despite significant advances in privacy and security, FL faces a fundamental trade-off between model expressiveness and computation efficiency. Research shows that to guarantee the ability to fit any target model, the rank of LoRA must meet a lower bound that scales with embedding size [113] as discussed in Section 3.4. However, implementing such high ranks in practice would incur prohibitive communication and computation costs, particularly challenging for resource-constrained devices [114, 115, 116]. To address this challenge, FedGBA [66] combines ensemble learning with rank-1 adaptations to enable highly efficient and expressive federated fine-tuning of FFM. Similarly, FFA-LoRA [110] reduces computational overhead by fixing randomly initialized non-zero matrices and fine-tuning only zero-initialized matrices.
Heterogeneity Handling. Another significant challenge in FL is managing heterogeneous data, devices and models across clients. FL models often suffer performance degradation when data across clients is non-independently and identically distributed (non-IID). SLoRA [117] addresses this through data-driven initialization, while FedLoRA [118] implements a stacking-based aggregation method enabling the aggregation of LoRA adapters with different ranks across heterogeneous clients. Additionally, pFedLoRA [119] introduces a homogeneous small adapter to facilitate federated clients’ heterogeneous local model training through iterative training for global-local knowledge exchange. Wagner et al. [120] introduce trust-weighted gradient aggregation schemes for on-device fine-tuning of large language models with limited data, utilizing LoRA to reduce communication and outperforming FedAvg and local tuning methods, particularly in diverse data scenarios. The approach effectively addresses data heterogeneity and scarcity in local datasets. Cho et al. [121] propose heterogeneous LoRA, which enables efficient federated fine-tuning of small foundation models on resource-limited devices by combining high and low-rank adaptations across clients, achieving optimal performance with minimal communication while preventing overfitting.
Personalization. Personalization in FL involves adapting the global model to individual clients through various techniques such as local fine-tuning, LoRA adaptation, or hybrid approaches. This process requires careful balancing of model performance, privacy protection, and system efficiency while considering each client’s unique data distribution and resource constraints. FedHLT [122] and FedLFC [123] exemplify this approach in Personalized Multilingual Modeling, combining low-rank adaptation with language family clustering to address FFM’s challenges of high communication costs and parameter interference, achieving improved performance with reduced overhead compared to baseline models. PER-PCS [124] is a framework that allows users to safely share and collaboratively assemble personalized LoRA pieces for large language models, achieving comparable performance to individual fine-tuning while significantly reducing computational costs and storage requirements.
The integration of LoRA with FFM represents a significant advancement in distributed machine learning, offering innovative solutions for privacy-preserving model adaptation while balancing efficiency, heterogeneity, and personalization requirements. Through various frameworks that combine privacy mechanisms, efficient parameter transmission, and adaptive techniques, LoRA have made FL more practical for resource-constrained environments while maintaining model expressiveness and security. As edge computing and IoT applications continue to evolve, the synergy between LoRA and FFM promises to transform distributed machine learning by enabling efficient, secure, and personalized model deployment at scale in privacy-sensitive domains.
4.5 LoRA for Long Sequence Modeling
The ability to process long sequences is crucial for many tasks across various domains handled by foundation models [125, 126, 127]. However, standard foundation models are typically constrained by their maximum context length due to the quadratic computational complexity of self-attention with respect to sequence length. To address this limitation, several LoRA-based have been proposed to extend the context window of foundation models.
Shifted Sparse Attention Method. LongLoRA [125] tackles the challenge by integrating position interpolation [128] with LoRA, enabling efficient fine-tuning of LLMs for longer contexts. Unlike standard LoRA applications, LongLoRA extends trainable low-rank adaptations to embedding and normalization layers in addition to the attention layers. A key innovation is the Shifted Sparse Attention (S2-Attn) mechanism, which approximates full attention during training by partitioning the input sequence into groups and applying attention within each group. To enhance information flow between groups, half of the attention heads are shifted by half the group size. This approach facilitates efficient training on extended sequences while preserving the original model architecture during inference.
Sink Fixed Attention Method. Building upon LongLoRA, SinkLoRA [126] introduces the Sink Fixed Attention (SF-Attn) mechanism to address specific limitations. SF-Attn combines a segmentation and reassembly algorithm with global attention focused on a limited number of ”sink attention tokens.” This method effectively redistributes attention scores, mitigating the overemphasis on initial tokens often observed in autoregressive models.
Another advancement, LongQLoRA[127], combines QLoRA[55] with position interpolation [128] and Shifted Short Attention. By quantizing the base model to 4-bit precision during fine-tuning, LongQLoRA enables context length extension with reduced computational resources compared to LongLoRA.
These LoRA-based techniques for long sequence modeling demonstrate significant potential in extending the context window of foundation models without incurring extensive computational overhead or necessitating full model fine-tuning.
4.6 LoRA Serving Systems
Efficient serving of multiple LoRA models is also essential. Recent advancements include improved GPU memory management [129], efficient batching techniques [130], CPU-assisted strategies to mitigate cold-start latency [131], and adaptation methods for resource-constrained personal devices [132].
S-LoRA [129] introduced a unified paging mechanism to manage both KV cache and LoRA weights in GPU memory, enabling concurrent serving of thousands of LoRA adapters. Punica [130] developed a custom CUDA kernel, Segmented Gather Matrix-Vector Multiplication (SGMV), facilitating efficient batching of requests across different LoRA models on a single GPU. CARASERVE [131] adopted a CPU-assisted strategy, initiating prefill computations for newly requested adapters during GPU loading to mitigate cold-start latency. Furthermore, CARASERVE introduced a rank-aware scheduling algorithm for optimized request routing in multi-GPU clusters. CA-LoRA [132] incorporated LoRA knowledge inheritance and model knowledge recovery mechanisms to maintain performance on personal devices. These innovations enhance the scalability and efficiency of multiple LoRA serving across different computing environments.
5 Application
The effectiveness and efficiency of LoRA in fine-tuning foundation models has led to its widespread adoption across various domains, including language processing, computer vision, speech recognition, multi-modal, code engineering, scientific discovery, recommender systems, graph learning, spatial-temporal forecasting, etc.
5.1 LoRA in Language Tasks
Language foundation models, such as LLaMA [4], RoBERTa [21] and DeBERTa [133], serve as important base models in LoRA study and have been extensively studied in different language tasks, including natural language understanding [47, 45, 48], question answering [36, 134], machine translation [49, 45], reasoning [56, 49, 75, 135] and natural language generation [55, 61, 62], which can also be seen from Table. This section explores its applications in specialized NLP domains.
Multilingual Language and Dialects Processing. LoRA enables efficient multilingual adaptation while preserving base models’ capabilities through minimal parameter updates. FedLFC [123] introduces a communication-efficient federated learning framework that uses LoRA for multilingual modeling while maintaining the base model’s weights and only updating lightweight LoRA parameters across language families. Similarly, FedHLT [122] employs LoRA with a hierarchical language tree strategy to enable efficient adaptation across dialects while minimizing communication overhead in federated learning settings. LAMPAT [136] leverages LoRA for unsupervised multilingual paraphrasing by applying adversarial training to generate diverse outputs while preserving semantic meaning across languages. HyperLoRA [137] develops a novel approach that generates dialect-specific LoRA adapters using linguistic feature vectors, enabling zero-shot adaptation to unseen English dialects without requiring dialectal training data.
Medical and Clinical Text Processing. Medical and clinical text processing faces unique challenges due to limited data availability and sensitive nature, especially in clinical datasets. LoRA has emerged as a promising solution to address these limitations. Ji et al. [138] utilized LoRA to enhance assertion detection in clinical settings, effectively improving classification accuracy with minimal data. Le et al. [139] also demonstrated that LoRA could effectively adapt pre-trained models for clinical NLP tasks, particularly in data-limited environments. Liu et al. [92] introduced a multi-task framework combining LoRA with MOE to address data imbalance in clinical multi-task scenarios. Shi et al. [140] utilized LoRA in their MedAdapter framework for effective test-time adaptation, thereby avoiding extensive computational resources or sharing data with third parties. Christophe et al. [141] conducted a comparative study with their proposed Med42 model, where LoRA fine-tuning surpassed traditional methods on key medical benchmarks like the USMLE.
5.2 LoRA in Computer Vision
LoRA has been effectively applied to various vision foundation models, like ViTs [145], DinoV2 [146], MAE [6], SAM [7], and Florence [147] enhancing their adaptability across multiple visual tasks including visual understanding, and visual generation.
5.2.1 Visual Understanding
Visual understanding contains a broad spectrum of tasks including domain adaptation, semantic segmentation, and content authenticity checking.
Domain Adaptation and Transfer Learning. Adapting foundation models trained on extensive natural image datasets to specialized domains such as medical imaging or satellite data often presents challenges due to limited domain-specific data and computational constraints. To address these challenges, several studies have explored the application of LoRA for efficient domain adaptation and transfer learning in various visual tasks.
ExPLoRA [148] extends pre-trained ViTs by incorporating LoRA modules within self-attention mechanisms, effectively modeling domain-specific style variations, such as those found in satellite imagery, in a low-dimensional space. Similarly, MeLo [149] demonstrates the efficacy of LoRA-based fine-tuning for medical imaging applications, including thoracic disease classification tasks. In addition, Kong et al. [150] apply LoRA to enhance the generalization of vision transformers for face forgery detection across different manipulation techniques and datasets.
A notable advancement in this direction is ConvLoRA [151], which extends the LoRA paradigm to convolutional neural networks for unsupervised domain adaptation in medical image. The architecture combines trainable low-rank decomposition matrices with adaptive batch normalization, establishing a more robust framework for domain transfer.
Semantic Segmentation. The adaptation of visual foundation models, particularly the Segment Anything Model (SAM) [7], has witnessed significant progress in semantic segmentation through LoRA integration. ConvLoRA [151] enhances SAM [7] for semantic segmentation in remote sensing, medical and agricultural images. Building upon this foundation, SAMed [152] demonstrates the effectivness of LoRA-based fine-tuning for multi-organ segmentation tasks. SurgicalSAM [153] applies similar techniques to the domain of robotic surgical instrument segmentation.
Content Authenticity Checking. Detecting synthesized content has become increasingly important with the advancement of generative models. CLIPMoLE [154] presents an architecture that adapts ViT blocks through a hybrid approach, combining shared and separate LoRA modules within an MoE framework. This design enables efficient fine-tuning of CLIP ViT models for transferable image detection across diverse generation techniques. Similarly, MoE-FFD [155] integrates LoRA with convolutional adapter modules while maintaining a frozen ViT backbone, specifically targeting face forgery detection.
5.2.2 Visual Generation
By training extra small networks, LoRA enables original pretrained models, such as diffusion models [158, 10], to be adapted to personalized styles and tasks or without retraining the entire foundation model.
Image Stylization. LoRA has emerged as a key technique for image stylization tasks due to its ability to rapidly adapt diffusion models to specific artistic styles while preserving the diverse generation capabilities of base model. [159] utilize LoRA to efficiently fine-tune Stable Diffusion [10] for comic-style transfer, specifically adapting it to the style of Calvin and Hobbes comics. Frenkel et al. [160] introduce B-LoRA, which leverages the architecture of Stable Diffusion XL [161] to implicitly separate style and content components of a single image. B-LoRA achieves style-content separation that enables various image stylization tasks, including image style transfer and text-based image stylization. Borse et al. [162] propose FouRA, which performs low-rank adaptation in the frequency domain rather than directly on model weights. FouRA addresses issues of distribution collapse and data copying that can occur when fine-tuning diffusion models with standard LoRA.
Multi-Concept Customization. Another common motivation for using LoRA in visual generation is to address the challenge of generating complex images involving multiple subjects or styles. Shah et al. [163] propose ZipLoRA, employing embedding-decomposed LoRA and gradient fusion to merge independently trained LoRAs without weight manipulation, preserving concept identities. Gu et al.[164] introduce Mix-of-Show, which utilizes LoRA switching and composition during the diffusion process. Similarly, Zhong et al.[165] investigate LoRA switching and composite approaches, alternating between or integrating LoRAs during de-noising to improve multi-element generation. Yang et al.[166] present LoRA-Composer, a training-free method integrating multiple LoRAs with concept injection and isolation constraints to mitigate issues like concept vanishing and confusion. Po et al. [167] propose orthogonal adaptation, which encourages customized weights of different concepts to be orthogonal, enabling efficient merging of independently fine-tuned models without compromising fidelity.
LoRA has also been applied to resolution-free generation tasks, ResAdapter by Cheng et al. [168] leverages LoRA to address the challenge of generating images at arbitrary resolutions while preserving their original style domain. Wang et al. [169] further advance this field with FiTv2, introducing enhanced model architecture and training strategies for resolution-free image generation.
Whether addressing multi-concept synthesis, high-resolution tasks, style transfer, or continual learning, LoRA provides a lightweight solution that enables models like Stable Diffusion and CLIP to maintain high performance while being adapted for new and complex tasks.
5.3 LoRA in Speech Recognition
LoRA has seen significant use in speech recognition tasks, particularly in efficiently fine-tuning foundation models like Wav2vec2 [8] and Whisper [170].
In fake audio detection, Wang et al. [171] applied LoRA to fine-tune the Wav2vec2 model [8], achieving comparable performance to full fine-tuning while reducing trainable parameters by a factor of 198 compared with the full fine-tuning. For multilingual automatic speech recognition (ASR), Xu et al. [172] proposed O-LoRA and O-AdaLoRA to adapt Whisper model [170] for low-resource languages like Uyghur and Tibetan. Similarly, Song et al [173] proposed LoRA-Whisper, that incorporates LoRA into Whisper, effectively mitigating language interference and facilitating the addition of new languages without degrading existing performance. Liu et al. [174] introduce LoRA-Whisper which extends LoRA to low-resource ASR by comparing it with bottleneck adapters for fine-tuning Whisper across seven low-resource languages.
5.4 LoRA in Code Engineering
In the field of code engineering, LoRA has emerged as a transformative approach for enhancing processes such as code review, repair, and generation tasks.
For code review and analysis, Lu et al. [175] introduced LLaMA-Reviewer, which fine-tunes LLaMA for review prediction, comment generation, and code refinement tasks with less than 1% trainable parameters. Silva et al. [176] developed RepairLLaMA, which employs a lightweight repair adapter for automated program repair, enabling effective deployment in resource-constrained environments.
For code generation and summarization, Kumar et al. [177] developed LoRA-based federated learning methods for code summarization, which preserves data privacy without direct access to source code. Cui et al. [178] introduced OriGen, which leverages code-to-code augmentation and self-reflection techniques for generating high-quality Register Transfer Level (RTL)222RTL is a design abstraction that describes digital circuits in terms of data flow between registers and their logical operations. code.
5.5 LoRA in Scientific Discovery
LoRA has been utilized across a wide range of scientific fields, including molecular tasks [179, 180, 96, 181, 182], and materials science [96].
In the field of protein analysis, Zeng et al.[179] developed PEFT-SP, a framework utilizing LoRA to fine-tune large protein language models (PLMs) like ESM-2 [183] for signal peptide prediction. This approach significantly improved performance, especially for rare peptide types, while mitigating overfitting and maintaining low computational costs. Similarly, Schmirler et al.[180] applied LoRA to PLMs, like like ProtT5 [184] and ESM-2, across various protein-related tasks, demonstrating accelerated training and improved downstream predictions while preventing catastrophic forgetting. In another molecular application, Schreiber [181] employed LoRA and its quantized version, QLoRA [55], in the ESMBind and QBind models for protein binding site and post-translational modification predictions. These models achieved enhanced generalization on unseen protein sequences without relying on structural data or multiple sequence alignments. Lv et al.[185] introduced ProLLaMA, demonstrating that a two-stage LoRA approach with different ranks could enable effective protein language learning while maintaining natural language capabilities.
In more general science area, Buehler et al. [96] developed X-LoRA, a framework that dynamically combines multiple specialized LoRA adapters. This approach demonstrated exceptional capability in solving inverse and forward tasks in protein mechanics and material design, illustrating LoRA’s potential for complex multidisciplinary applications.
5.6 LoRA in Recommender Systems
In recommender systems, LoRA efficiently fine-tunes LLMs for CTR prediction and sequential recommendation tasks.
For CTR prediction, Yang et al. [186] proposed MLoRA, implementing domain-specific low-rank matrices to capture inter-domain variations and enhance personalization, successfully deployed at Alibaba. Zhu et al. [187] introduced RecLoRA, which replaces single LoRA matrices with a meta-LoRA structure, using soft routing to select personalized matrix combinations for each user’s evolving interests.
For sequential recommendation, Qin et al. [188] introduced ATFLRec which integrates audio and text data by independently optimizing LoRA modules for each modality. Zheng et al. [189] developed LLM-TRSR for text-rich recommendation scenarios, leveraging LoRA’s efficient fine-tuning to handle large-scale text data while maintaining real-time capabilities. Kong et al. [190] developed iLoRA that employs an MoE framework, creating specialized LoRA modules for distinct user behavior patterns, dynamically engaging relevant experts based on individual interaction sequences. Recently, Ji et al.[191] proposed GenRec, a pure text-based LLM that uses item names as IDs and leverages LoRA to fine-tune LLaMA, showing superior performance on large-scale datasets where LLMs can directly learn collaborative information from text representations.
5.7 LoRA in Graph Learning
Recently, researchers have also explored the applications of LoRA on non-Euclidean data (i.e., graphs), fine-tuning graph neural networks (GNN) [192] to adapt to new graphs or structure updates of existing graphs.
For cross-domain graph neural network adaption, Yang et al.[193] introduced GraphLoRA, which constructs a small trainable GNN alongside the pre-trained one to bridge structural and feature distribution gaps between different graphs. By applying low-rank decomposition to the tunable GNN parameters and incorporating graph structure-based regularization, GraphLora can effectively adapt pre-trained GNNs to diverse graph domains by fine-tuning only 20% of the parameters.
For dynamic knowledge graph learning, Liu et al. [194] developed IncLoRA, which adapts entity and relation embeddings to continuous graph updates. IncLoRA groups new knowledge embeddings into explicit LoRA layers based on their distance to preserved graphs and allocates rank scales adaptively using graph structural properties.
5.8 LoRA in Spatial-Temporal Forecasting
Multivariate time series data is prevalent in real-world scenarios such as transportation, weather forecasting, and economics [195]. Recent studies have explored using LoRA to address specific challenges in this field, such as node-specific adaptation and multi-channel modeling and out-of-domain prediction.
For node-specific adaptation, Ruan et al. [196] proposed ST-LoRA, which implements node-adaptive LoRA layers to add extra learnable parameters for each node. By incorporating residual structure among LoRA layers to avoid overparameterization, ST-LoRA effectively captures distinct patterns and dynamics of different nodes in transportation datasets.
For multi-channel modeling, Nie et al. [197] introduced C-LoRA, which balances between channel-dependent and channel-independent strategies. C-LoRA parameterizes each channel with a low-rank factorized adapter to form identity-aware embeddings, then inputs these to a globally shared predictor for modeling cross-channel dependencies.
For out-of-domain prediction, Gupta et al.[198] analyzed the effectiveness of LoRA-based fine-tuning across leading time series foundation models like Lag-Llama[199], MOIRAI [200] and Chronos [201], showing improved vital sign of sepsis patients in intensive care units forecasting while reducing computational costs. Ren et al.[202] introduced TPLLM, which injects trainable rank-decomposition matrices into GPT-2 Transformer blocks for traffic forecasting, effectively adapting the model to process spatial-temporal representations with limited historical traffic data.
5.9 LoRA in Multi-Modal
Multimodal Foundation Models (MFMs) combine different data modalities including text, audio, images, video, etc within a shared representational space, enabling cross-modal reasoning and understanding. LoRA enhances these models by optimizing training efficiency while strengthening inter-modal alignments. Language-vision and language-audio learning represent two main domains where LoRA has been extensively applied to adapt MFMs.
5.9.1 Language-vision Learning
In Language-vision learning tasks, LoRA has been applied particularly in enhancing vision-language capabilities and customizing diffusion models.
Language-Vision Model Based Adaptation. Sung et al. [203] adapted Vision-Language Model (VLM) by fine-tuning CLIP-BART for visual question answering and image captioning. Ji et al. [204] enhanced the adversarial robustness of VLMs through clustering-based LoRA for cross-modal retrieval. In addition, Zong et al. [205] proposed MoVA to efficiently route and fuse multiple vision foundation models (CLIP [15], DINOv2 [146], SAM [7]) through a coarse-to-fine adapter mechanism.
Diffusion Model Based Customization. Various approaches have leveraged LoRA to efficiently adapt Stable Diffusion (SD) models. For concept customization, Kumari et al. [206] introduced Custom Diffusion with LoRA in domain adapter layers for few-shot concept learning, while Li et al. [207] developed SELMA to train and merge multiple skill-specific LoRA experts without interference. Lu et al. [208] advanced concept manipulation through LoRA modules with cross-attention refinement for selective concept erasure. Additional applications include StitchDiffusion [209] for 360-degree panorama generation, Chinese garden image synthesis [210], DreamSync [211] for improved generation faithfulness, Block-wise LoRA [212] for fine-grained control, and AnimateDiff [213] for motion adaptation.
In vision-language model adaptation, LoRA has enhanced cross-modal understanding through techniques like CLIP-BART fine-tuning [203], clustering-based approaches [204], and multi-model fusion [205]. For diffusion model customization, LoRA has enabled efficient concept learning [206], selective manipulation [208], and specialized generation tasks ranging from panorama creation [209] to motion adaptation [213].
5.9.2 Language-Audio Learning
In Language-Audio Learning tasks, LoRA has emerged as a pivotal technique for addressing the fundamental challenge of modality gap between audio and text representations in foundation models. The applications primarily fall into two categories: speech recognition and audio content understanding and generation.
Speech Recognition. Fathullah et al. [214] introduced a direct audio embedding method where LoRA adapts the self-attention layers specifically for audio-text alignment, enabling large language models to process audio inputs effectively. Fathullah et al. [214] demonstrated that LLMs can perform multilingual ASR by prepending audio embeddings to text tokens, adapting only critical model parameters through LoRA while achieving competitive performance against specialized ASR systems. Yusuf et al. [215] proposed speculative speech recognition by combining an RNN-Transducer-based ASR system [216] with an audio-prefixed language model adapted using LoRA, allowing the model to generate speculative predictions. Palaskar et al. [217] further introduced FLoRA with modality-specific adapters that can be selectively engaged, providing a robust framework for multimodal integration.
Speech Content Understanding and Generation. Liu et al. [218] proposed LOAE for automated audio captioning (AAC), which combines CED-based audio encoding [219] with LLaMA for text decoding, bridged by Q-Former [220], and optimized through LoRA-based adaptation. Qin et al. [188] introduced ATFLRec, a multimodal recommender system that leverages separate LoRA modules to efficiently fine-tune audio and text modalities in LLMs, demonstrating superior performance through their fusion-based approach.
These developments in language-audio learning have established several key principles for applying LoRA in multi-modal contexts: (1) the importance of modality-specific adaptation paths [214, 217], (2) the effectiveness of selective parameter updating for cross-modal alignment [215, 218], and (3) the value of maintaining distinct representational spaces through specialized LoRA modules before fusion [217, 188].
6 Challenges and Discussion
While LoRA has demonstrated remarkable success in efficiently adapting foundation models across diverse domains, several critical challenges and opportunities further further investigation.
Theoretical Understanding. Current theoretical frameworks primarily focus on simplified settings or specific architectures [72], leaving gaps in our understanding of LoRA’s behavior in more complex scenarios. For instance, the interaction between LoRA adaptations and the pretrained model’s knowledge is not fully understood, particularly in terms of how LoRA preserves useful features while modifying task-specific ones. Developing more comprehensive theoretical frameworks that can explain LoRA’s effectiveness across different architectures and tasks would be valuable for guiding future improvements.
Architectural Design Principles. Current LoRA implementations often rely on empirical observations rather than systematic design methodologies [221, 17]. Critical questions persist regarding optimal adapter placement strategies, rank determination across network depths, and the geometric properties of adaptation spaces. Recent explorations into non-Euclidean geometries, such as hyperbolic spaces [222], suggest potential advantages for capturing hierarchical relationships in model adaptations. A unified framework for analyzing these design choices could significantly advance our understanding of the parameter efficiency versus model capacity trade-off.
Computational Efficiency. The scalability of LoRA becomes increasingly crucial as LLMs continue to expand. Managing concurrent adaptations while handling variable-length sequences presents significant challenges in memory utilization and computational overhead. The dynamic management of adapter modules and KV cache tensors can lead to memory fragmentation and increased I/O costs [129]. Advanced serving architectures and optimization techniques are essential for maintaining low latency in production environments, particularly when dealing with multiple concurrent adaptation requests.
Robustness and Verification. The deployment of LoRA-adapted models in critical applications necessitates robust verification methods. Current research inadequately addresses model behavior under distribution shifts and adversarial conditions [223]. Developing rigorous uncertainty quantification methods and formal verification techniques becomes paramount, especially for high-stakes applications in healthcare and autonomous systems where model reliability directly impacts human safety.
Privacy and Security. As LoRA becomes more widely adopted, particularly in federated learning settings, privacy and security considerations become increasingly important. This includes protecting sensitive information in LoRA adaptations, preventing unauthorized access or manipulation of adapted models, and ensuring that LoRA updates do not inadvertently leak private information [110, 224]. Research into privacy-preserving LoRA adaptation techniques and secure methods for sharing and combining LoRA modules is needed.
Future research directions should focus on: establishing comprehensive theoretical frameworks that unify various LoRA design aspects; developing automated architecture search methods for optimal adapter configurations; creating efficient serving infrastructures for large-scale deployments; implementing standardized evaluation protocols for reliability assessment; and integrating advanced privacy-preserving mechanisms into the adaptation process.
Moreover, as novel architectures emerge, such as Mamba [225], investigating LoRA’s applicability to these new paradigms becomes crucial. The integration of LoRA into edge computing and real-time systems [226] presents additional challenges that intersect with hardware optimization and system design, necessitating interdisciplinary research efforts.
7 Conclusion
In this survey, we have presented a systematic analysis of LoRA, examining its theoretical underpinnings, technical advances, and diverse applications in adapting foundation models. The extensive adoption of LoRA across diverse domains—from natural language processing and computer vision to speech recognition and scientific computing—highlight its versatility and effectiveness. Its success in maintaining model performance while significantly reducing computational and storage requirements has made it particularly valuable for resource-constrained environments and specialized domain adaptations.
Despite these achievements, several critical challenges persist. The theoretical framework underlying LoRA’s effectiveness requires further development, particularly in understanding the interaction between low-rank adaptations and model capabilities. Additionally, questions regarding scalability, robustness, and secure deployment in production environments present ongoing research opportunities.